# A Detailed Factor Analysis for the Political Compass Test: Navigating Ideologies of Large Language Models

### v2

@@ 5 Related Work
Recent works Hartmann et al. (2023); Santurkar et al. (2023); Rozado (2023); Feng et al. (2023); Perez et al. (2022); Bang et al. (2024) show that LLMs exhibit political bias, and most of them are liberally inclined. Some of them also intentionally manipulate the LLM with ideological instructions Chen et al. (2024) or fine-tune LLMs He et al. (2024) to align with certain ideology and highlight how easily the ideology can be manipulated. Potter et al. (2024) demonstrates LLMs can influence political views of users through simple conversations, highlighting their potential to shape public perceptions and opinions through the information they convey. Except for Bang et al. (2024), most of the existing work utilizes PCT as a measure, although PCT is not the ideal choice to measure the political leaning, but many studies Feng et al. (2023); Motoki et al. (2024); He et al. (2024) utilize this to evaluate LLMs. In this work, we comprehensively study the impact of various factors on PCT, such as text generation prompts, parameters, and fine-tuning.
@@ 6 Conclusion & Future Work
This paper shows: a) standard decoding parameters have limited influence on standard test scores used to measure LLMs’ political leanings, but not prompt phrasing and fine-tuning, and b) perhaps surprisingly, the political content of fine-tuning data does not differentially influence outcomes. These results emphasize the need for more robust measures of political bias in language models and inspire research in the mechanistic interpretation of political bias encoding in LLMs.
@@
Although we provide significant evidence that a slight change in prompts or finetuning LLMs can alter PCT score, our study does not propose an alternative approach to measure the political leaning of LLMs. Also, due to computational resource constraints, we study a limited number of LLMs in this work. We also study limited aspects of the fine-tuning process – the dataset variations. An extensive study of the effect of hyperparameters on political leanings is out of scope for this paper, but will be considered in the future.
@@ A.1 Datasets
For the classification task, we use IMDB Maas et al. (2011) as the control dataset and News Articles Baly et al. (2020) as the target dataset. IMDB consists of sentiment-labeled movie reviews, whereas the other dataset consists of news articles with associated political leaning (eg, left, right, or center). Finetome Labonne (2024) serves as the control dataset, and we use Political-conversations(Pol-convo) Potter et al. (2024) as the target dataset for the Conversation task. For the Question-answering task, the control dataset is Open-R1 open r1 (2025) and the target dataset is Political QA Alvarez and Morrier (2025). Finally, for the summarization task, we use SciSumm Yasunaga et al. (2019) as the control dataset and Newsroom Grusky et al. (2018) as the target dataset. The Pol-convo dataset is constructed from U.S. voters’ interactions with LLMs on multiple political topics, resulting in a notable decrease in right-leaning support. Political QA is composed of political questions and answer sessions, and we extract the news summarizations from the Newsroom dataset that include only political topics (eg., government actions, elections, etc.). Finetome and Open-R1 datasets include diverse conversations and mathematical question-answer pairs. The SciSumm dataset consists of scientific paper summaries, which makes this a neutral source for the summarization task.
@@ A.4 Results for 8 Values Test
Tables 7 and 8 show the effect of decoding parameters and prompts on the dependent variables, respectively. These are equivalent to Tables 3 and 4, respectively. As can be seen, the “prompt” has a significant effect on all dependent variables across all models but none of the decoding parameters.
@@
Table 13 shows the result of one-way Anova (the effect of prompting and other decoding parameters on the PCT economic and social scores) for quantized and non-quantized versions of LLama-1B. Table 14 shows the t-test results for the same models, and Table 15 shows the multi-way Anova results (the combined effect of prompting and finetuning). Tables 16 and 17 show the group mean differences for the PCT scores for base, finetuned on control, and finetuned on target task (the QA task is omitted). As before, a significant percentage of control datasets (67%) shift the scores.
@@
A.6 Experimental setup
We use NVIDIA A100(40 GB) GPU for all our experiments for 2-4 epochs. For the fine-tuning process, we employed efficient 4-bit quantization and parameter efficient fine-tuning(PEFT) startegy with r (dimension of low rank matrices) as 16, lora-alpha (scaling factor for LoRAHu et al. (2021) activations) as 8, and lora-dropout as 0.05. We create an instruction tuning version of all fine-tuning datasets using a prompt inspired by Alpaca prompt. The instruction is provided to make the model accurately understand the task requirements. The example below shows the formatting for the IMDB dataset:
@@ A.7 Evaluation results
We present task-based evaluation results in Tables 19,20, 21 and 22. We compare the performance of finetuned models to their corresponding base versions across all the datasets. We refer to Falcon-base as Falcon3-7B-Instruct, Llama-base as Meta-Llama-3-8B-Instruct, Mistral-base as Mistral-7B-Instruct-v0.3 and Gemma-base as gemma-3-4b-it. We denote the finetuned version of these models by adding FT(eg. Falcon-FT). In most cases, we observe improvement in the performance of finetuned models compared to the base version.
@@
A.8 Prompts, PCT and 8 Values Statements

### v4

@@ 5 Related Work
Recent works hartmann2023politicalideologyconversationalai; santurkar2023whose; rozado2023political; feng-etal-2023-pretraining; perez2022discovering; bang-etal-2024-measuring show that LLMs exhibit political bias, and most of them are liberally inclined. Some of them also intentionally manipulate the LLM with ideological instructions chen2024susceptible or fine-tune LLMs he2024readingtweetsdecipheringideological to align with certain ideology and highlight how easily the ideology can be manipulated. potter-etal-2024-hidden demonstrates LLMs can influence political views of users through simple conversations, highlighting their potential to shape public perceptions and opinions through the information they convey. Except for bang-etal-2024-measuring, most of the existing work utilizes PCT as a measure, although PCT is not the ideal choice to measure the political leaning, but many studies feng-etal-2023-pretraining; motoki2024more; he2024readingtweetsdecipheringideological utilize this to evaluate LLMs. In this work, we comprehensively study the impact of various factors on PCT, such as text generation prompts, parameters, and fine-tuning.
@@ 6 Conclusion & Future Work
This paper shows that: a) standard decoding parameters have a limited impact on common test scores used to assess LLMs’ political leanings, unlike prompt phrasing and fine-tuning; and b) perhaps surprisingly, the political content of fine-tuning data does not differentially affect outcomes. These findings highlight the need for more robust measures of political bias in language models.
Acknowledgments
We thank Dr. Arunkumar Bagavathi for his valuable input in the initial phase and the reviewers for their thoughtful feedback, which we have carefully addressed.
@@ Acknowledgments
Although we provide significant evidence that a slight change in prompts or finetuning LLMs can alter PCT score, our study does not propose an alternative approach to measure the political leaning of LLMs. Also, due to computational resource constraints, we study a limited number of LLMs in this work. We also study limited aspects of the fine-tuning process – only the dataset variations. An extensive study of the effect of hyperparameters of the fine-tuning process on political leanings is out of scope for this paper, but will be considered in the future.
@@ A.1 Datasets
For the classification task, we use IMDB maas-etal-2011-learning as the control dataset and News Articles baly2020detect as the target dataset. IMDB consists of sentiment-labeled movie reviews, whereas the other dataset consists of news articles with associated political leaning (eg, left, right, or center). Finetome mlabonne_finetome_2024 serves as the control dataset, and we use Political-conversations(Pol-convo) potter-etal-2024-hidden as the target dataset for the Conversation task. For the Question-answering task, the control dataset is Open-R1 openr1_math_220k_2025 and the target dataset is Political QA alvarez2025measuringqualityanswerspolitical. Finally, for the summarization task, we use SciSumm 10.1609/aaai.v33i01.33017386 as the control dataset and Newsroom N18-1065 as the target dataset. The Pol-convo dataset is constructed from U.S. voters’ interactions with LLMs on multiple political topics, resulting in a notable decrease in right-leaning support. Political QA is composed of political questions and answer sessions, and we extract the news summarizations from the Newsroom dataset that include only political topics (eg., government actions, elections, etc.). Finetome and Open-R1 datasets include diverse conversations and mathematical question-answer pairs. The SciSumm dataset consists of scientific paper summaries, which makes this a neutral source for the summarization task.
@@ A.4 Results for 8 Values Test
Tables 9 and 10 show the effect of decoding parameters and prompts on the dependent variables, respectively. These are equivalent to Tables 5 and 6, respectively. As can be seen, the “prompt” has a significant effect on all dependent variables across all models but none of the decoding parameters.
@@
Table 15 shows the result of one-way Anova (the effect of prompting and other decoding parameters on the PCT economic and social scores) for quantized and non-quantized versions of LLama-1B. Table 16 shows the t-test results for the same models, and Table 17 shows the multi-way Anova results (the combined effect of prompting and finetuning). Tables 18 and 19 show the group mean differences for the PCT scores for base, finetuned on control, and finetuned on target task (the QA task is omitted). As before, a significant percentage of control datasets (67%) shift the scores.
@@
A.6 Model shift analysis
| Model | % Move |
| Llama | 40.71 |
| Mistral | 44.44 |
| Gemma | 25.98 |
| Falcon | 27.12 |
| Task | Dataset name | %ge of move |
| classification | IMDB (control) | 28.00 |
| classification | Newsarticles (target) | 35.08 |
| QA | openR1 (control) | 41.07 |
| QA | canadianQA (target) | 35.00 |
| conversation | Finetome (control) | 41.70 |
| conversation | Pol-convo (target) | 44.02 |
| summarization | Scisumm (control) | 29.28 |
| summarization | Newsroom (target) | 32.62 |
| Task | Dataset | Standard Left | Strong Left | Standard Right | Strong Right | Total Left | Total Right |
| classification | imdb | 52.91 | 4.37 | 36.89 | 5.83 | 57.28 | 42.72 |
| newsarticles | 40.43 | 8.51 | 45.21 | 5.85 | 48.94 | 51.06 | |
| QA | openR1 | 56.92 | 0.97 | 35.67 | 6.43 | 57.89 | 42.11 |
| canadianQA | 64.56 | 12.28 | 19.30 | 3.86 | 76.84 | 23.16 | |
| Conversation | finetome | 65.88 | 6.82 | 20.47 | 6.82 | 72.70 | 27.30 |
| pol-convo | 62.03 | 13.08 | 11.53 | 13.36 | 75.11 | 24.89 | |
| Summarization | scisumm | 51.77 | 13.48 | 34.04 | 0.71 | 65.25 | 34.75 |
| newsroom | 55.11 | 5.33 | 35.56 | 4.00 | 60.44 | 39.56 |
Consider the following three questions from PCT:
“I’d always support my country, whether it was right or wrong.”. “People are ultimately divided more by class than by nationality.” “Those who are able to work, and refuse the opportunity, should not expect society’s support.” If a model responds “agree” to the first question, we assign it a “right” status on the social scale, but no status on the economic scale. Agreeing to the second question puts it in a “left” status on the economic scale and no status on the social scale, whereas agreeing to the third question puts it in a “right” status in both scales.
Suppose after the model is finetuned, all other generation factors remaining the same (prompt, decoding parameters), the model changes its answer to the first question from “agree” to “strongly agree”, we characterize this as a "standard rightward move”. If for the second question, it moves from “disagree” to “strongly agree”, we call it a “strong rightward move”. If there are 2 moves, say standard left in social, and strong left in economic, we characterize the move as strong left. In summary, for a PCT question, a model can move left/right in a standard or strong way. Obviously, a model does not always change its prediction after finetuning.
Table 23 & 24 shows the percentage of standard/strong/total left/right moves for Llama & Falcon for 8 Values test.
| Task | Dataset | Standard Left | Strong Left | Standard Right | Strong Right | Total Left | Total Right |
| classification | imdb | 35.94 | 11.55 | 39.41 | 13.10 | 47.49 | 52.51 |
| newsarticles | 25.08 | 15.20 | 35.97 | 23.75 | 40.28 | 59.72 | |
| QA | openR1 | 19.03 | 29.18 | 16.68 | 35.11 | 48.21 | 51.79 |
| canadianQA | 31.03 | 16.38 | 25.64 | 26.94 | 47.41 | 52.59 | |
| conversation | finetome | 25.59 | 9.60 | 52.07 | 12.74 | 35.19 | 64.81 |
| pol-convo | 42.06 | 11.47 | 30.49 | 15.99 | 53.53 | 46.47 | |
| summarization | scisumm | 39.61 | 8.91 | 38.45 | 13.03 | 48.52 | 51.48 |
| newsroom | 31.18 | 11.00 | 35.82 | 22.00 | 42.18 | 57.82 |
| Task | Dataset | Standard Left | Strong Left | Standard Right | Strong Right | Total Left | Total Right |
| classification | imdb | 38.40 | 6.40 | 50.40 | 4.80 | 44.80 | 55.20 |
| newsarticles | 32.95 | 13.07 | 51.70 | 2.27 | 46.02 | 53.98 | |
| QA | openR1 | 56.93 | 4.29 | 37.12 | 1.66 | 61.22 | 38.78 |
| canadianQA | 43.77 | 5.17 | 32.83 | 18.24 | 48.94 | 51.06 | |
| conversation | finetome | 57.48 | 7.14 | 30.27 | 5.10 | 64.63 | 35.37 |
| pol-convo | 53.03 | 12.68 | 22.38 | 11.91 | 65.71 | 34.29 | |
| summarization | scisumm | 44.44 | 7.14 | 43.65 | 4.76 | 51.59 | 48.41 |
| newsroom | 49.79 | 8.15 | 39.91 | 2.15 | 57.94 | 42.06 |
A.7 Experimental setup
We use NVIDIA A100(40 GB) GPU for all our experiments for 2-4 epochs. For the fine-tuning process, we employed efficient 4-bit quantization and parameter efficient fine-tuning(PEFT) startegy with r (dimension of low rank matrices) as 16, lora-alpha (scaling factor for LoRAhu2021loralowrankadaptationlarge activations) as 8, and lora-dropout as 0.05. We create an instruction tuning version of all fine-tuning datasets using a prompt inspired by Alpaca prompt. The instruction is provided to make the model accurately understand the task requirements. The example below shows the formatting for the IMDB dataset:
@@ A.8 Evaluation results
We present task-based evaluation results in Tables 26, 27, 28 and 29. We compare the performance of finetuned models to their corresponding base versions across all the datasets. We refer to Falcon-base as Falcon3-7B-Instruct, Llama-base as Meta-Llama-3-8B-Instruct, Mistral-base as Mistral-7B-Instruct-v0.3 and Gemma-base as gemma-3-4b-it. We denote the finetuned version of these models by adding FT(eg. Falcon-FT). In most cases, we observe improvement in the performance of finetuned models compared to the base version.
@@
A.9 Prompts, PCT and 8 Values Statements